# MINIMIZATION OF NONSMOOTH WEAKLY CONVEX FUNCTION OVER PROX-REGULAR SET FOR ROBUST LOW-RANK MATRIX RECOVERY

### v1

@@ 5 Concluding Remarks
In this paper, we proposed a prox-regular-type low-rank constrained optimization model, incorporating a weakly convex loss function, for robust low-rank matrix recovery. For the proposed model, we presented an optimization algorithm, for minimization of a weakly convex function over a prox-regular set, with guaranteed convergence in terms of stationary point. Our numerical experiments demonstrate that the proposed model with a weakly convex loss function, called SCAD, dramatically improves estimation performance of robust low-rank matrix recovery under severe outliers. Finally, we defer further numerical evaluations and complete proofs, including the extensions discussed in Remark 3.5, to an extended manuscript in preparation.

### v2

@@ 5 Concluding Remarks
In this paper, we proposed a prox-regular-type low-rank constrained optimization model, incorporating a weakly convex loss function, for robust low-rank matrix recovery. For the proposed model, we presented an optimization algorithm, for minimization of a weakly convex function over a prox-regular set, with guaranteed convergence in terms of stationary point. Our numerical experiments demonstrate that the proposed model with a weakly convex loss function, called SCAD, dramatically improves estimation performance of robust low-rank matrix recovery under severe outliers. Finally, we defer further numerical evaluations and complete proofs to an extended manuscript in preparation.